# M3D: Manifold-based Domain Adaptation with Dynamic Distribution for Non-Deep Transfer Learning in Cross-subject and Cross-session EEG-based Emotion Recognition

### v2

@@ III-B Dynamic Distribution Alignment
Given the absence of label information for the target domain during the training phase, we cannot directly obtain $f(z_{t}^{(c)})$ in (12). To tackle this issue, we approximate $f(z_{t}^{(c)})$ using the class-conditional probability distribution [34]. Specifically, the process begins by training an initial weak classifier, such as a SVM or KNN, on the source data. This initial classifier, $f$, is subsequently utilized to infer pseudo-labels for the target data, serving as a provisional substitute for the unavailable true labels of the target domain. To enhance the reliability and accuracy of the pseudo-labels, an iterative refinement strategy is employed to improve the decodability $f$. This involves using the outcomes of the initial classifier to iteratively train subsequent classifiers, thereby progressively refining the pseudo-labels. This iterative process aims to converge towards more accurate pseudo-labels, thus improving the model’s ability to generalize from the source to the target domain effectively. Further details on the specific iterative refinement process and how it contributes to classifier learning will be discussed in Section III-C.
In (12), the adaptive factor $\mu$ is not a constant value predetermined by prior knowledge. Instead, it is dynamically learned based on the underlying data distribution. Specifically, we use the A-distance measurement [35] as a metric to estimate the marginal distribution difference, denoted as $d_{A}$, between the source and target domains, as
@@
where $\epsilon\left(h\right)$ denotes the hinge loss, which is derived from a binary classifier trained specifically to distinguish between samples from the source and target domains. For the conditional distribution difference, based on the provided label information for source data and the estimated pseudo-label information of target data, we conduct the A-distance measurement within each $c$ class ($c\in[1,\dots,C]$) as
@@
In the process of dynamic distribution alignment, the adaptive factor $\mu$ is recalculated at each iteration, underscoring the iterative and responsive nature of this adaptation strategy. This re-calibration is essential to ensuring that the adaptation process remains attuned to the evolving similarities and differences between the source and target domain feature distributions as they are represented in the manifold feature space.
III-C Classifier Learning
@@
Here, $\iota$ is the iteration number, $y_{i}$ is the groundtruth, and $f(z_{i})$ is the predicted classification result. $\|\cdot\|^{2}_{H_{K}}$ refers to the L2 norm in the reproducing kernel Hilbert space. The dynamic distribution adaptation term, $\overline{D}_{f}\left(T_{S},T_{T}\right)$, is obtained in (12). Further, a Laplacian regularization term, $R_{f}\left(T_{S},T_{T}\right)$, is incorporated to leverage the inherent geometric similarities among neighboring points in the manifold $\mathbb{G}$ [30]. Here, the classifier $f$ can be initialized using a commonly employed machine learning classifier, followed by adaptive optimization through an iterative learning process.
@@
$\begin{split}R_{f}\left(T_{S},T_{T}\right)=\sum^{n+m}_{i,j=1}W_{ij}\left(f\left(z_{i}\right)-f\left(z_{j}\right)\right)^{2}\\ =\sum^{n+m}_{i,j=1}f\left(z_{i}\right)L_{ij}f\left(z_{j}\right),\\ \end{split}$ (23)
@@
1:- Data matrix $X=\left[X_{s},X_{t}\right]$, source labels $y_{s}$, subspace dimension $d$, regularization parameters $\lambda,\beta,\rho$ and the iteration number $\iota$;
2:Classifier $f(\cdot)$;
3:# Manifold Feature Transformation
4:Learn the manifold feature transformation kernel $G$ and obtain the manifold feature representation $z$ using (8);
5:Train a weak classifier based on the source data $D_{S}$, and get initialized soft labels $\hat{y}_{t}^{0}$ for the target data $D_{T}$;
6:Construct the geodesic line kernel $K$ using the transformed features $z_{s}=z_{1:n}$ and $z_{t}=z_{n+1:n+m}$;
7:for $\iota=1$ to $l$ do
8:# Dynamic Distribution Alignment
9:     Calculate the adaptive factor $\mu$ using (15), and compute $M_{0}$ and $M_{c}$ using (21) and (22);
10:# Classifier Learning
11:     Solve the equation using (27) to compute $\beta^{*}$ and obtain the classifier $f$ using (26);
12:     Update the soft labels for the target data $\hat{y}_{t}^{\iota}=f(z_{t})$;
13:end for
14:Return the classifier $f$;
15:# Ensemble Learning
16:Perform ensemble learning based on the obtained soft labels in the loop $\{\hat{y}_{t}^{\iota}\}_{\iota=1}^{l}$.
III-D Ensemble Learning
@@
IV Experimental Results
IV-A Emotional Databases
The model performance is carefully validated on two well-known publicly available databases: SEED [11] , SEED-IV [37]. The SEED database includes EEG emotion data collected from 15 participants. Each participant was exposed to a total of 15 movie clips designed to evoke three types of emotions (positive, negative, and neutral). The SEED-IV database also includes EEG emotion data from 15 participants, exposed to a total of 24 movie clips aimed at eliciting four types of emotions (happy, sad, neutral, and fear). For both SEED and SEED-IV databases, participants underwent three sessions spread across separate days, with a one-week break between sessions. The EEG signals were recorded using the ESI Neuroscan system with 62 channels.
The EEG preprocessing followed established procedures in the literature. The process began with the downsampling of EEG signals to 200 Hz, followed by the removal of artifacts such as the Electrooculogram (EOG) and Electromyography (EMG). A bandpass filter with a 0.3-50 Hz range was then applied to improve the signal quality. Following this, each trial was segmented into multiple 1-second data samples. To extract emotional related information, DE features were extracted across five frequency bands: Delta (1-3 Hz), Theta (4-7 Hz), Alpha (8-13 Hz), Beta (14-30 Hz), and Gamma (31-50 Hz). Thus, for each 1-second data sample, a total of 310 features (5 frequency bands $\times$ 62 channels) were characterized, which serves as the input for the model.
To further assess the generalizability of the proposed M3D model beyond controlled experimental conditions, we extend model validation to a real-world clinical EEG database obtained from Hospital Universiti Sains Malaysia (HUSM). This dataset consists of EEG recordings from 27 healthy individuals (mean age: 38.28±15.64 years) and 29 individuals diagnosed with MDD (Major Depressive Disorder) (mean age: 40.33±12.86 years). Participant classification adhered to the international diagnostic criteria outlined in the Diagnostic and Statistical Manual-IV (DSM-IV). This dataset used different devices and settings compared to the SEED and SEED-IV databases, with EEG recordings collected using the Brain Master Discovery amplifier, 19 electrodes, and a sampling rate of 256 Hz. Preprocessing involved bandpass filtering (0.5-70 Hz) and notch filtering at 50 Hz to eliminate power-line noise. The experimental protocol included EEG recordings under both open-eye (EO) and closed-eye (EC) conditions, each lasting 5 minutes. During the EO condition, participants were instructed to remain relaxed while minimizing eye movements to maintain data quality.
IV-B Experimental Protocols
  • Cross-subject single-session leave-one-subject-out cross-validation. The data samples from N subjects (SEED/SEED-IV: N=14) in the first session are used as the source domain, and the data samples from the remaining 1 subject in the same session are used as the target domain. This procedure repeats N+1 times (SEED/SEED-IV: N=14), ensuring that each participant is treated as the target domain at least once.
  • Cross-subject cross-session leave-one-subject-out cross-validation. The data samples from N subjects (SEED/SEED-IV: N=14) across all three sessions are used as the source domain, and the data samples from the remaining 1 subject across all three sessions are used as the target domain. This procedure also repeats N+1 times (SEED/SEED-IV: N=14), ensuring that each participant is treated as the target domain at least once.
In the present study, we adopt six widely recognized machine-learning classifiers to generate the initial classification results ($\hat{y}_{t}^{0}$) as mentioned in Section III-C. The adopted machine-learning classifiers include KNN[38], SVM[39], DT[40], Adaboost classifier[41], GNB classifier[42], and Bagging classifier[43]. In the SVM classifier, the radial basis function (RBF) kernel is utilized. In ensemble learning, the application of three similarity matrix functions (CTS, SRS, and ASRS) combined with three hierarchical agglomerative clustering algorithms (SL, CL, and AL) yields a total of 9 possible combinations. Here, the combination that achieves the highest accuracy in predicting target domain labels will be chosen as the optimal outcome. Notably, all the best results are achieved using the DT classifier and the CTS-SL approach. For performance evaluation, a comprehensive analysis of the model’s efficacy is conducted using the accuracy and F1-score metrics.
IV-C Experimental Results on Cross-Subject Single-Session
The experimental results on SEED and SEED-IV database under the cross-subject single-session cross-validation are reported in Table I. For the SEED database, the best performance is obtained when DT is adopted as the initial classifier, where the corresponding accuracy and F1-score are 84.57% and 76.11%. We compare the obtained experimental results with the existing literature, as shown in Table II. The comparison results show that the proposed non-deep transfer learning model (M3D) demonstrates superior performance. In comparison to the best results obtained by the traditional machine learning methods, our model shows a performance enhancement of 8.26%. Furthermore, it performs comparably or even surpasses some deep learning methods, such as PGCN [44], ADA [15], TANN [45] and DG-DANN [46].
TABLE I: Cross-subject single-session leave-one-subject-out cross-validation on SEED and SEED-IV with different initial classifiers.
@@ IV-C Experimental Results on Cross-Subject Single-Session
F1-score 75.77 75.84 76.11 74.49 74.49 74.49
@@
TABLE II: Performance comparison on SEED using cross-subject single-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
@@
SVM[13] 58.18/13.85 TKL[47] 63.54/15.47
TCA[13] 64.00/14.66 SCA[46] 66.33/10.60
SA[48] 69.00/10.89 KPCA[13] 69.02/09.25
DICA[46] 69.40/07.80 GFK[47] 71.31/14.09
@@
MLP[13] 61.01/12.38 DNN[13] 61.01/12.38
DCORAL[49] 62.14/07.98 GELM[50] 70.80/00.74
DAN[49] 73.54/06.09 DDC[49] 74.34/09.05
R2G-STLT[51] 77.96/06.38 DANN[13] 79.79/13.14
MS-MDA[49] 79.67/08.01 DGCNN[52] 79.95/09.02
MMD [15] 80.88/10.10 SimNet[53] 81.58/05.11
GECNN[54] 82.46/10.83 Bi-DANN[47] 83.28/09.60
DGGN[50] 83.84/10.26 Saliency[55] 84.11/02.90
BiDANN-S[47] 84.14/06.87 R2G-STNN[56] 84.16/07.63
DG-DANN[46] 84.30/08.32 TANN[45] 84.41/08.75
ADA[15] 84.47/10.65 PGCN[44] 84.59/08.68
BiHDM[48] 85.40/07.53 ST-SCGNN[57] 85.90/04.90
@@
For the SEED-IV database, the best accuracy is obtained when DT is used as the initial classifier, where the corresponding accuracy and F1-score are 60.94% and 50.49%, respectively. The performance comparison results with the existing methods are reported in Table III. Compared to the best results reported in the traditional machine learning methods, the proposed M3D demonstrates a slight performance advantage of 0.72%. Furthermore, it performs comparably or even surpasses some deep learning methods, including ScalingNet [51], ROADAN [58], WGAN-GP [59] and MS-ADA [49].
TABLE III: Performance comparison on SEED-IV using cross-subject single-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
@@
TCA[60] 37.01/10.47 GFK[60] 44.04/09.31
GFHF[60] 49.29/07.60 JDA[60] 54.57/05.47
GAKT[60] 58.49/06.23 MIDA[60] 60.22/08.69
@@
DCORAL[49] 40.50/10.05 DAN [49] 42.03/05.66
DDC[49] 44.06/06.01 JD-IRT[61] 52.12/15.88
RGNN[62] 53.49/10.57 JAGP[62] 54.37/19.49
DANN[63] 54.63/08.03 GECNN[55] 55.01/10.29
MS-MDA[64] 59.34/05.48 MS-ADA[49] 59.29/13.65
WGAN-GP[59] 60.60/15.76 ROADAN[58] 60.75/10.40
ScalingNet[51] 60.75/10.40 MFA-LR[62] 64.96/14.04
@@
IV-D Experimental Results on Cross-Subject Cross-Session
To evaluate the efficiency and stability of the proposed M3D in managing variations in EEG data collected from the same subject at different times, we also conduct a model validation under the cross-subject cross-session leave-one-subject-out cross-validation protocol. The experimental results are reported in Table IV, when different initial classifiers are used. The best model performance of SEED database is achieved when DT is adopted as the initial classifier, where the accuracy and F1-score are 77.05% and 66.46%, respectively. The performance results against the existing methods are compared in Table V. It shows, the proposed model demonstrates an accuracy improvement of 5.18% in comparison to traditional machine learning methods. Further, it exhibits performance that is comparable, to some extent, with deep learning methods, showing the effectiveness of the proposed M3D in managing variations in data collected from the same subject across different time periods.
TABLE IV: Cross-subject cross-session leave-one-subject-out cross-validation on SEED and SEED-IV with different initial classifiers
@@ IV-D Experimental Results on Cross-Subject Cross-Session
TABLE V: Performance comparison on SEED using cross-subject cross-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
@@
MetaEmotionNet[70] 77.50/08.80 DGCNN[72] 79.95/09.02
DAN [73] 81.04/05.32 DANN[74] 81.08/05.88
DCORAL[75] 81.97/05.16 DDC[76] 82.17/04.96
@@
For the SEED-IV database, the performance across various initial classifiers is relatively consistent, with the highest accuracy reaching 64.90% when DT is employed as the initial classifier. Table VI reports the performance comparison results against the existing machine learning and deep learning methods. Compared to the best results reported in various literature on machine learning methods, the proposed M3D shows a performance improvement of 0.46%. Furthermore, it performs comparably or even outperforms some deep learning methods, including MetaEmotionNet [70], MS-ADA [49] and DAN [13].
TABLE VI: Performance comparison on SEED-IV using cross-subject cross-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
@@
KNN[65] 40.83/07.28 GFK[28] 45.89/08.27
CORAL[67] 49.44/09.09 RF[68] 50.98/09.20
SVM[13] 51.78/12.85 KPCA[13] 51.76/12.89
TPT[13] 52.43/14.43 Adaboost[69] 53.44/09.12
TCA[13] 53.97/08.05 SA[48] 64.44/09.47
@@
DCORAL[49] 40.50/10.05 DAN[49] 42.03/05.66
DDC[49] 44.06/06.01 ACRNN[70] 49.20/09.20
DNN[13] 49.35/09.74 MLP[13] 49.35/09.74
MMResLSTM[70] 51.10/11.40 DGCNN[52] 52.82/09.23
STRNN[70] 53.20/07.40 3D-CNN[70] 54.10/10.90
CDCN[70] 54.50/13.10 DANN[13] 54.63/08.03
A-LSTM[71] 55.03/09.28 TSception[70] 56.20/09.50
STFFNN[70] 56.70/07.30 MS-MDA[49] 57.92/10.12
DAN [13] 58.87/08.13 MS-ADA[49] 59.29/13.65
MetaEmotionNet[70] 61.20/08.30 Bi-DANN[47] 65.59/10.39

### v3

@@ 3.2 Dynamic Distribution Alignment
Given the absence of label information for the target domain during the training phase, we cannot directly obtain $f(z_{t}^{(c)})$ in (12). To tackle this issue, we approximate $f(z_{t}^{(c)})$ using the class-conditional probability distribution [31]. Specifically, the process begins by training an initial weak classifier, such as a SVM or KNN, on the source data. This initial classifier, $f$, is subsequently utilized to infer pseudo-labels for the target data, serving as a provisional substitute for the unavailable true labels of the target domain. To enhance the reliability and accuracy of the pseudo-labels, an iterative refinement strategy is employed to improve the decodability $f$. This involves using the outcomes of the initial classifier to iteratively train subsequent classifiers, thereby progressively refining the pseudo-labels. This iterative process aims to converge towards more accurate pseudo-labels, thus improving the model’s ability to generalize from the source to the target domain effectively. Further details on the specific iterative refinement process and how it contributes to classifier learning will be discussed in Section 3.3.
In (12), the adaptive factor $\mu$ is not a constant value predetermined by prior knowledge. Instead, it is dynamically learned based on the underlying data distribution. This parameter comprehensively incorporates both the marginal distribution difference and conditional distribution difference between domains. Specifically, we use the A-distance measurement [32] as a metric to estimate the marginal distribution difference, denoted as $d_{A}$, between the source and target domains, as
@@
where $\epsilon\left(h\right)$ denotes the hinge loss from a binary classifier distinguishing between samples from the source and target domains. For the conditional distribution difference, based on the provided label information for source data and the estimated pseudo-label information of target data, we conduct the A-distance measurement within each $c$ class ($c\in[1,\dots,C]$) as
@@
In the process of dynamic distribution alignment, the adaptive factor $\mu$ is recalculated at each iteration, underscoring the iterative and responsive nature of this adaptation strategy. This re-calibration is essential to ensure that the adaptation process remains attuned to the evolving similarities and differences between the source and target domain feature distributions as they are represented in the manifold feature space. More detailed descriptions of $\mu$ are presented in the Supplementary Materials Appendix A.
3.3 Classifier Learning
@@
Here, $\iota$ is the iteration number, $y_{i}$ is the groundtruth, and $f(z_{i})$ is the predicted classification result. $\|\cdot\|^{2}_{H_{K}}$ refers to the L2 norm in the reproducing kernel Hilbert space. The dynamic distribution adaptation term, $\overline{D}_{f}\left(T_{S},T_{T}\right)$, is obtained in (12). Further, a Laplacian regularization term, $R_{f}\left(T_{S},T_{T}\right)$, is incorporated to leverage the inherent geometric similarities among neighboring points in the manifold $\mathbb{G}$ [27]. Here, the classifier $f$ can be initialized using a commonly employed machine learning classifier, followed by adaptive optimization through an iterative learning process. $\eta,\lambda$ and $\rho$ are regularization parameters.
@@
$\begin{split}R_{f}\left(T_{S},T_{T}\right)&=\sum^{n+m}_{i,j=1}W_{ij}\left(f\left(z_{i}\right)-f\left(z_{j}\right)\right)^{2}\\ &=\sum^{n+m}_{i,j=1}f\left(z_{i}\right)L_{ij}f\left(z_{j}\right),\\ \end{split}$ (23)
@@
1:
2:- Data matrix $X=\left[X_{s},X_{t}\right]$, source labels $y_{s}$, subspace dimension $d$, regularization parameters $\eta,\lambda,\rho$ and the iteration number $\iota$;
3:Classifier $f(\cdot)$;
4:# Manifold Feature Transformation
5:Learn the manifold feature transformation kernel $G$ and obtain the manifold feature representation $z$ using (8);
6:Train a weak classifier based on the source data $D_{S}$, and get initialized soft labels $\hat{y}_{t}^{0}$ for the target data $D_{T}$;
7:Construct the geodesic line kernel $K$ using the transformed features $z_{s}=z_{1:n}$ and $z_{t}=z_{n+1:n+m}$;
8:for $\iota=1$ to $l$ do
9:# Dynamic Distribution Alignment
10:     Calculate the adaptive factor $\mu$ using (15), and compute $M_{0}$ and $M_{c}$ using (21) and (22);
11:# Classifier Learning
12:     Solve the equation using (27) to compute $\beta^{*}$ and obtain the classifier $f$ using (26);
13:     Update the soft labels for the target data $\hat{y}_{t}^{\iota}=f(z_{t})$;
14:end for
15:Return the classifier $f$;
16:# Ensemble Learning
17:Perform ensemble learning based on the obtained soft labels in the loop $\{\hat{y}_{t}^{\iota}\}_{\iota=1}^{l}$.
3.4 Ensemble Learning
@@
Table 1: Cross-subject single-session LOSO CV on SEED, SEED-IV and SEED-V with different initial classifiers.
@@
Sensitivity 82.57 80.82 76.04 81.08 81.08 81.08
Specificity 85.29 86.25 88.78 85.42 85.42 85.42
Precision 71.07 73.48 76.94 70.80 70.80 70.80
F1-score 76.39 76.97 76.49 75.59 75.59 75.59
AUROC 88.37 88.36 88.36 88.04 88.04 88.04
NPV 91.79 90.51 88.27 91.19 91.19 91.19
@@
AUROC 71.65 72.20 73.68 72.32 71.89 71.69
NPV 79.31 80.75 88.02 79.63 85.03 79.31
SEED-V Accuracy 64.39 64.41 65.25 63.05 63.44 63.04
Sensitivity 62.30 62.30 57.53 60.48 62.30 62.30
Specificity 64.55 64.57 66.16 63.33 63.52 63.10
Precision 11.58 11.59 16.60 15.40 11.29 11.18
F1-score 19.54 19.54 25.76 24.55 19.12 18.96
AUROC 77.87 77.80 78.46 76.92 77.28 77.10
NPV 95.83 95.83 93.01 93.56 95.76 95.74
4 Experimental Results
4.1 Emotional Databases
The model performance is carefully validated on three well-known publicly available databases: SEED [12], SEED-IV [34] and SEED-V[35]. The SEED database includes EEG emotion data collected from 15 participants. Each participant was exposed to a total of 15 movie clips designed to evoke three types of emotions (positive, negative, and neutral). The SEED-IV database also includes EEG emotion data from 15 participants, exposed to a total of 24 movie clips aimed at eliciting four types of emotions (happy, sad, neutral, and fear). In the SEED-V database, a total of 5 emotions (happiness, sadness, fear, disgust, and neutral) are elicited using 15 movie clips, and the simultaneous EEG signals under the 5 emotions are recorded from 16 subjects (6 males and 10 females). For both the SEED, SEED-IV, and SEED-V databases, participants undergo three sessions spread across separate days, with a one-week break between sessions. The EEG signals are recorded using the ESI Neuroscan system with 62 channels.
The EEG preprocessing follows established procedures in the literature. This approach aligns with the standard preprocessing practices in emotion-related EEG research and ensures consistency across the widely used SEED, SEED-IV, and SEED-V datasets in the literature. The process begins with the downsampling of EEG signals to 200 Hz, followed by the removal of artifacts such as the Electrooculogram (EOG) and Electromyography (EMG). Downsampling to 200 Hz effectively reduces computational complexity while preserving the temporal resolution necessary for emotion recognition [12, 34, 35]. A bandpass filter with a 0.3-50 Hz range is then applied to improve the signal quality, as the majority of emotion-related electroencephalography activities occur below 50 Hz. Following this, each trial was segmented into multiple 1-second data samples. To extract emotional related information, DE features were extracted across five frequency bands: Delta (1-3 Hz), Theta (4-7 Hz), Alpha (8-13 Hz), Beta (14-30 Hz), and Gamma (31-50 Hz). Thus, for each 1-second data sample, a total of 310 features (5 frequency bands $\times$ 62 channels) were characterized, which serves as the input for the model. To explore EEG signal inter-subject differences in SEED and SEED-IV datasets, we perform a comprehensive statistical analysis of input EEG features via qualitative visualizations and quantitative comparisons. More detailed descriptions and results are presented in Supplementary Materials Appendix B.
To further assess the generalizability of the proposed M3D model beyond controlled experimental conditions, we extend model validation to a real-world clinical EEG database obtained from Hospital Universiti Sains Malaysia (HUSM) [36]. This dataset consists of EEG recordings from 27 healthy individuals (mean age: 38.28±15.64 years) and 29 individuals diagnosed with MDD (Major Depressive Disorder) (mean age: 40.33±12.86 years). Participant classification adhered to the international diagnostic criteria outlined in the Diagnostic and Statistical Manual-IV (DSM-IV). This dataset used different devices and settings compared to the SEED and SEED-IV databases, with EEG recordings collected using the Brain Master Discovery amplifier, 19 electrodes, and a sampling rate of 256 Hz. Preprocessing involved bandpass filtering (0.5-70 Hz) and notch filtering at 50 Hz to eliminate power-line noise. This preprocessing pipeline is widely used and validated in the literature [36] as a reliable approach for enhancing data quality and processing efficiency without compromising informative neural features. The experimental protocol included EEG recordings under both open-eye (EO) and closed-eye (EC) conditions, each lasting 5 minutes. During the EO condition, participants were instructed to remain relaxed while minimizing eye movements to maintain data quality. Similar to the SEED, SEED-IV and SEED-V databases, this database also extracts DE features across five frequency bands.
4.2 Experimental Protocols
  • Cross-subject single-session leave-one-subject-out cross-validation (Cross-subject single-session LOSO CV). The data samples from N subjects (SEED/SEED-IV: N=14, SEED-V: N=15) in the first session are used as the source domain, and the data samples from the remaining 1 subject in the same session are used as the target domain. This procedure repeats N+1 times (SEED/SEED-IV: N=14, SEED-V: N=15), ensuring that each participant is treated as the target domain at least once.
  • Cross-subject cross-session leave-one-subject-out cross-validation (Cross-subject cross-session LOSO CV). The data samples from N subjects (SEED/SEED-IV: N=14, SEED-V: N=15) across all three sessions are used as the source domain, and the data samples from the remaining 1 subject across all three sessions are used as the target domain. This procedure also repeats N+1 times (SEED/SEED-IV: N=14, SEED-V: N=15), ensuring that each participant is treated as the target domain at least once.
In the present study, we adopt six widely recognized machine-learning classifiers to generate the initial classification results ($\hat{y}_{t}^{0}$) as mentioned in Section 3.3. The adopted machine-learning classifiers include KNN[37], SVM[38], DT[39], Adaboost classifier[40], GNB classifier[41], and Bagging classifier[42]. In the SVM classifier, the radial basis function (RBF) kernel is utilized. In ensemble learning, the application of three similarity matrix functions (CTS, SRS, and ASRS) combined with three hierarchical agglomerative clustering algorithms (SL, CL, and AL) yields a total of 9 possible combinations. Here, the combination that achieves the highest accuracy in predicting target domain labels will be chosen as the optimal outcome. Notably, all the best results are achieved using the DT classifier and the CTS-SL approach. For performance evaluation, a comprehensive analysis of the model’s effectiveness is conducted using multiple metrics, including accuracy, sensitivity, specificity, precision, F1-score, Area Under the Receiver Operating Characteristic Curve (AUROC) and Negative Predictive Value (NPV) metrics. We set the manifold feature dimension $d=128$ in all datasets except for the MDD database (where $d=100$). The regularization parameters are set as $\eta=0.1$, $\rho=1$ and $\lambda=0.4$.
@@ 4.2 Experimental Protocols
Table 2: Performance comparison on SEED using cross-subject single-session LOSO CV. The second-best traditional machine learning method result is underlined.
@@
SVM[14] 58.18/13.85 TCA[14] 64.00/14.66
KPCA[14] 69.02/09.25 SCA[43] 66.33/10.60
DICA[43] 69.41/07.79 KLIEP[44] 45.71/17.76
ULSIF[44] 51.18/13.57 TKL[44] 63.54/15.47
SA[44] 69.00/10.89 GFK[44] 71.31/14.09
@@
MLP[14] 61.01/12.38 DANN[14] 79.79/13.14
DAN[14] 83.81/08.56 MMD [16] 80.88/10.10
ADA[16] 84.47/10.65 DG-DANN[43] 84.30/08.32
DGCNN[44] 79.95/09.02 Bi-DANN[44] 83.28/09.60
BiDANN-S[44] 84.14/06.87 DCORAL[45] 62.14/07.98
DDC[45] 74.34/09.05 MS-MDA[45] 79.67/08.01
SOGNN[46] 76.00/06.92 R2G-STLT[46] 77.96/06.38
A-LSTM[47] 72.18/10.85 TANN[47] 84.41/08.75
GECNN[48] 82.46/10.83 Saliency[48] 84.11/02.90
P-GCNN[48] 84.35/10.28 ST-SCGNN[48] 85.90/04.90
@@
Table 3: Performance comparison on SEED-IV using cross-subject single-session LOSO CV. The second-best traditional machine learning method result is underlined.
@@
GFK[15] 44.04/09.31 GFHF[15] 49.29/07.60
JDA[15] 54.57/05.47 GAKT[15] 58.49/06.23
MIDA[15] 60.22/08.69 KLIEP[47] 31.46/09.20
ULSIF[47] 32.99/11.05 STM[47] 39.39/12.40
SVM[47] 37.99/12.52 TCA[47] 56.56/13.77
@@
ScalingNet[46] 60.75/10.40 GECNN[48] 55.01/10.29
DANN[47] 47.59/10.01 DGCNN[47] 52.82/09.23
A-LSTM[47] 55.03/09.28 DAN[47] 58.87/08.13
RGNN[49] 53.49/10.57 JAGP[49] 54.37/09.49
MFA-LR[49] 64.96/14.04 DDC[50] 37.41/06.36
DCORAL[50] 37.43/03.08 MS-MDA[50] 59.34/05.48
@@
Table 4: Performance comparison on SEED-V using cross-subject single-session LOSO CV. The second-best traditional machine learning method result is underlined. Models reproduced by the authors are indicated with an asterisk (*).
Methods $P_{acc}$ Methods $P_{acc}$
Traditional machine learning methods
TCA*[18] 39.32/13.18 SA*[23] 33.45/09.92
GFK* [25] 36.44/16.14 SVM*[38] 49.39/18.13
Adaboost*[40] 50.01/21.71 KNN*[51] 33.14/13.62
KPCA*[52] 34.41/19.21 CORAL*[53] 51.10/21.12
Deep learning methods
DDC[54] 28.70/09.40 JDA[54] 29.74/06.90
DAN[54] 41.24/12.12 JD-IRT[54] 60.17/11.06
M3D 65.25/07.87
Table 5: Cross-subject cross-session LOSO CV on SEED, SEED-IV and SEED-V with different initial classifiers
@@
AUROC 76.24 76.24 76.46 76.43 76.35 76.35
NPV 82.51 82.51 84.99 86.43 82.57 82.57
SEED-V Accuracy 56.37 56.61 59.15 56.61 56.36 56.37
Sensitivity 54.97 54.97 58.65 54.70 54.70 54.70
Specificity 56.64 56.93 59.24 56.98 56.69 56.70
Precision 19.85 19.96 20.60 20.00 19.89 19.89
F1-score 29.17 29.29 30.49 29.29 29.17 29.17
AUROC 72.76 72.92 74.70 72.92 72.77 72.76
NPV 86.55 86.61 88.82 86.48 86.42 86.42
Table 6: Performance comparison on SEED using cross-subject cross-session LOSO CV. The second-best traditional machine learning method result is underlined.
@@
MetaEmotionNet[55] 77.50/08.80 A-LSTM[56] 72.18/10.85
DGCNN[56] 79.95/09.02 EEG-SWTNS[56] 80.07/10.75
@@
Table 7: Performance comparison on SEED-IV using cross-subject cross-session LOSO CV. The second-best traditional machine learning method result is underlined.
@@
KPCA[14] 51.76/12.89 TPT[14] 52.43/14.43
KNN[15] 40.83/07.28 Adaboost[15] 53.44/09.12
CORAL[15] 49.44/09.09 RF[55] 34.70/06.60
SVM[55] 41.10/07.40 TCA[56] 56.56/13.77
SA[56] 64.44/09.46 GFK[56] 64.38/11.41
@@
DANN[14] 54.63/08.03 DAN [14] 58.87/08.13
MLP[55] 50.70/06.30 STFFNN[55] 56.70/07.30
ACRNN[55] 49.20/09.20 MMResLSTM[55] 51.10/11.40
TSception[55] 56.20/09.50 STRNN[55] 53.20/07.40
3D-CNN[55] 54.10/10.90 CDCN[55] 54.50/13.10
MetaEmotionNet[55] 61.20/08.30 A-LSTM[56] 55.03/09.28
DGCNN[56] 52.82/09.23 EEG-SWTNS[56] 66.72/10.19